the sale price of smart Tv is 40% lesser tha the sale price of smartphone . If the regular price of smartphone is Php 12,550.00 is on sale at 20% off . What is the sale price of the smart TV?


1
Expert's answer
2021-12-14T05:31:15-0500

Price of smartphone on sale at 20% off "=80/100\times 12550= Php 10040"

The sale price of smart Tv is 40% lesser than the sale price of smartphone,

The 40% is equivalent to "40\% \times 10040=Php4016"

The sale of smart Tv "=10040-4016=Php6024"